The area of a square is given by 9z^2+48z+64. Express the width of the square in terms of Z and express the perimeter in terms of Z
So far I have (3x+8)(3x+8), not to sure where to go from here.

Answer by rothauserc(4718)

9z^2 +48z +64 = x^2, where x is the length of a side of the square
factor the polynomial
(3z+8)^2 = x^2
take square root of both sides of =
x = 3z +8
therefore side of square x = 3z +8
now the perimeter(P) of the square is
P = 4*(3z+8) = 12z +32